Suppose a construction company is trying to decide whether to buy a new nail gun. The table below shows the hypothetical costs for the nail gun and the amount the gun will save the company each year. Assume the gun will last forever. In each case, determine the highest interest rate the company should pay for a loan that makes purchase of the nail gun possible.



1
Expert's answer
2021-03-29T12:04:50-0400

assume the table is as shown below


rate of interest = r

Cost = Savings per year/ rate of interest

Case A:

Highest rate of interest the company would pay:

r = 100/1000 = 10%

Case B:

r = 200/1000 = 20%

Case C:

r = 300/1000 = 0.3(100%)=30%